# UPS: Efficiently Building Foundation Models for PDE Solving via Cross-Modal Adaptation

### v2

@@ 3.2 Unified Architecture
Large-scale transformer models have demonstrated success in various domains, spanning natural language [e.g., 39], vision [e.g., 40], and audio processing [e.g., 41]. In this work, we explore the potential of transformers for PDE solving. We break down the UPS architecture into three parts: an embedding network that transforms the unified representation into sequential features; the model body, consisting of the pretrained LLM layers; and a predictor that generates the prediction (Figure 2).
@@
We use FNO due to its strong empirical performance [4, 10] and its ability to extract resolution-invariant features. As we consider maximum two-dimensional PDEs in this work, we use a series of 2D FNO layers with $l$ channels to obtain PDE features in $\mathbb{R}^{l\times n^{d}}$. Then, to map the FNO output to the LLM’s embedding dimension, we apply a pointwise convolution with input channel $n^{d}$, output channel $e$, kernel size $1$, stride $1$. This yields the desired sequential features $h_{\mathrm{PDE}}\in\mathbb{R}^{l\times e}$.
Since UPS is intended to handle diverse data from various generating sources, we leverage the PDE’s metadata in addition to the input dynamics. The motivation is that LLMs can use the textual information to better understand the context and characteristics of different PDEs. To implement this, we specify the metadata in the form “[PDE family][coefficients]" which is embedded into sequential features $h_{\mathrm{meta}}$ using the LLM’s tokenizer and text embedding layer. We then concatenate the meta features and the PDE features to obtain $h_{\mathrm{mix}}:=[h_{\mathrm{meta}},h_{\mathrm{PDE}}]$. Finally, we apply positional encoding and layer normalization to $h_{\mathrm{mix}}$. This will be the input to the subsequent transformer layers.
@@
Linear Predictor Finally, we define a prediction head to transform the hidden state of the LLM body $\hat{h}$ to the predicted next step of the input $\hat{\bm{u}}_{t+1}^{s}(\bm{x})\in\mathbb{R}^{N\times n^{d}}$ (we predict all the physical quantities in the set $V$). This is achieved by averaging over the sequence dimension of $\hat{h}$ to get shape $\mathbb{R}^{e}$, applying a linear layer to map it to $\mathbb{R}^{Nn^{d}}$, and reshaping the results to obtain the final prediction $\hat{\bm{u}}_{t+1}^{s}(x)$.
@@ 4 Full Workflow and Training
Stage 1: Embedding Pretraining Intuitively, there is a modality gap between text data used to train general-purpose LLMs and PDEs. Previous work has also shown that directly fine-tuning pretrained LLMs on non-text inputs can result in suboptimal performance [42]. To address this, Shen et al. [16] introduced ORCA, which performs distribution matching before fine-tuning to enable cross-modal adaptation. That is, given a randomly initialized embedding network, we first pretrain it to minimize the distribution distance between the embedding network’s output—in our case $h_{\mathrm{mix}}$—and the text embeddings of a external reference NLP dataset, which we denote as $h_{\mathrm{LM}}$. This process makes the cross-modal distribution resemble the text distribution that the LLM is pretrained on.
@@
$\begin{split}{\mathcal{L}}_{\mathrm{align}}&=\|\mu_{\mathcal{D}_{\mathrm{h_{mix}}}}-\mu_{\mathcal{D}_{\mathrm{h_{LM}}}}\|_{L_{2}}\\ &=\mathbb{E}_{\mathcal{D}_{\mathrm{h_{mix}}}}[k(a,a^{\prime})]-2\mathbb{E}_{\mathcal{D}_{\mathrm{h_{mix}}},\mathcal{D}_{\mathrm{h_{LM}}}}[k(a,b)]-\mathbb{E}_{\mathcal{D}_{\mathrm{h_{LM}}}}[k(b,b^{\prime})]\end{split}$ (2)
@@ 5 Experiments
Data We train and evaluate our method using PDEBench [10]. For training, we combine 7 datasets from different PDE families: Burgers Equation (1D), Advection (1D), Diffusion-Sportion (1D), Shallow-Water (2D), compressible Navier-Stokes (1D and 2D), and incompressible Navier-Stokes (2D). We explicitly hold out two families, 1D and 2D Diffusion-Reaction, to evaluate the generalization ability of UPS. The dataset details can be found in Appendix B. We use the scale-independent normalized root mean squared error (nRMSE) as the evaluation metric, defined as follows:
@@
We preprocess all the PDEs by normalizing each dataset along the channel dimension to ensure the scale of $\bm{u}^{s}_{t}(\bm{x})$ across datasets is similar. The model’s prediction is denormalized to compute nRMSE.
Baselines We compare against two sets of baselines: (i) single-family models trained on individual PDE datasets, including the widely used U-Net [43], FNO [33], the improved version FFNO [44], the transformer-based GNOT [45] and OFormer [46], as well as the cross-modal ORCA [16]; (ii) unified models trained on multiple datasets, including MPP [13], DPOT [14], and a unified FNO trained using data transformed by our unified representation scheme. We note that MPP and DPOT focus on 2D PDEs and are pretrained on 2D Navier-Stokes, Shallow-Water, and Diffusion-Reaction from PDEBench. Subramanian et al. [12] is not included as a baseline because its models are pretrained on different PDE families (e.g., Poisson’s and Helmholtz equations) not present in PDEBench.
Implementation Details As noted in Section 3, UPS is compatible with any pretrained LLM. We present our main results using RoBERTa [21] and study other backbones in ablation studies (Table 4). We set the embedding FNO channel $l$ to 32. Since the resolution of the 2D datasets in PDEBench is 128, we set the model resolution $n$ to 128 and downsample datasets with higher resolutions. All of our experiments can be run on a single NVIDIA A6000 GPU. See Appendix C for training details.
Table 1: nRMSEs (lower is better) for in-distribution PDEBench families, with baseline results taken from Takamoto et al. [10], Shen et al. [16], McCabe et al. [13], Hao et al. [14]. ‘-’ means that the result is not available. On all datasets, UPS with RoBERTa-Base (UPS-B) achieves the lowest nRMSEs among all smaller models and UPS with RoBERTa-Large (UPS-L) achieves the lowest nRMSEs among all large models. Numbers are bolded for each model size group.
@@ 5.1 Achieving State-of-the-Art Results on PDEBench with Compute Efficiency
Since MPP and DPOT focus on 2D problems and use a different set of pretraining datasets from ours, we train a 2D-only UPS on all 2D datasets in PDEBench to provide a more direct comparison. The results are shown in Appendix Table 7. Notably, while 2D UPS is still trained with less data (since the other methods use additional datasets like PDEArena [11]), our method ranks first on 4 of 8 datasets, outperforming DPOT on 5 of 8 datasets and outperforming MPP on 3 of 4 datasets.
@@
In addition to prediction errors in Table 1, we visualize some of the UPS outputs in Appendix E and show that it is indeed able to capture the key features and dynamics of different PDE families.
@@ 5.2 Generalizing to Unseen PDEs with Data Efficiency
Unseen PDE Families As mentioned earlier, we hold out the Diffusion-Reaction equations from developing UPS. We first directly evaluate UPS on these two tasks and report the zero-shot transfer performance. Then, we study few-shot transfer by randomly sampling $k\in\{10,100\}$ trajectories from the training sets of the held-out tasks and use them to fine-tune UPS. Lastly, we report the fine-tuning results with the full training dataset. The results are shown in Table 2. As the number of adaptation samples increases, the prediction error decreases. Notably, the 100-shot result of UPS on
Table 2: Zero- and few-shot transfer performance of UPS on unseen PDE families and coefficients. Our few-shot results are competitive with baselines trained with more data.
Unseen PDE Families Unseen Coeff
# Samples Model 1D Diff-React 2D Diff-React Burgers $\nu=1.0$
UPS-B 0.0557 1.0593 0.0566
0 FNO 0.1839 1.2 1.0342
ORCA 0.1818 1.0812 1.6316
UPS-B 0.0107 0.3327 0.0134
10 FNO 0.1698 0.8193 0.67
ORCA 0.1004 0.5376 0.4829
UPS-B 0.0034 0.2508 0.0022
100 FNO 0.0037 0.1869 0.0123
ORCA 0.0051 0.1362 0.027
UPS-B 0.0003 0.041 0.0008
9K (Full) FNO 0.0014 0.12 0.0031
ORCA 0.0034 0.082 0.012
1D datasets is better than the baselines trained on 9,000 data. That is, we use 90x less data to match the performance of single-family operators. This makes UPS useful for low-resource PDE problems where data collection is costly and training models from scratch is challenging. On 2D Diffusion-Reaction, we are slightly worse than pretrained MPP (0.0292) and DPOT (0.0106) since this dataset is considered as in-distrubution for MPP and DPOT.
Unseen Coefficients UPS also generalizes to PDEs in the same families as the training data but with different coefficients. We verify this by adapting UPS to Burgers Equation with $\nu=1.0$ (the model is trained on $\nu=0.001$). The last column in Table 2 shows that while our zero-shot performance is already competitive, the performance after further adaptation outperforms all considered baselines.
Table 3: UPS with resolution 128 has an nRMSE of 0.0033. We directly test it on higher resolutions.
@@
Advection nRMSE 0.0057 0.0064 0.0068
@@
As shown in Table 3, although the nRMSEs for the Advection Equation slightly increase compared to the nRMSE for the training resolution, they still outperform all baselines in Table 1. Since the numbers are similar across columns, UPS generalizes to higher resolutions in a zero-shot manner.
@@ 5.3 Ablation Studies
Table 4: Results for the ablation studies. For each set of experiments, only the specified settings are different; all the other hyperparameters and training configurations are the same. Overall, the full UPS-B workflow (first row for every study) most effectively leverages the pretrained knowledge of LLMs and obtains the best results.
@@
S3 Concatenation 0.0027 0.0399 0.0009 0.0056 0.0016 0.0153 0.0931
Cross-Attention 0.003 0.0420 0.0009 0.0065 0.0023 0.0189 0.1082
No Metadata 0.0122 0.0453 0.001 0.0091 0.0026 0.0238 0.1171
@@
S2: Cross-Modal Alignment We also test the importance of the two objectives used in stage 1, i.e., alignment loss with MMD, and task loss with nRMSE. We study three settings: (i) using only ${\mathcal{L}}_{\mathrm{align}}$ for stage 1 as in Shen et al. [16]; (ii) using only ${\mathcal{L}}_{\mathrm{task}}$ for stage 1; and (iii) removing stage 1 from our workflow entirely. As shown in Table 4, while removing any objective reduces the performance across all datasets, removing the task loss has a more significant negative effect. Meanwhile, removing the entire stage of embedding pretraining hurts prediction accuracy. This shows that simply fine-tuning the LLM without considering the modality gap or learning to extract PDE features is ineffective.
@@
S3: Incorporating Text-Form Metadata UPS leverages the PDE’s metadata by combining its text embeddings with the learned PDE embeddings. To study whether incorporating metadata is helpful and identify an optimal approach, we compare our workflow with two alternatives: (i) we do not use metadata, so $h_{\mathrm{mix}}:=h_{\mathrm{PDE}}$; (ii) we use metadata, but instead of concatenating features from two modalities, we apply a cross-attention mechanism: $h_{\mathrm{mix}}:=\mathrm{softmax}(\frac{QK^{T}}{\sqrt{e}})V$, where $Q=W_{Q}h_{\mathrm{PDE}}$, $K=W_{K}h_{\mathrm{meta}}$, and $V=W_{V}h_{\mathrm{meta}}$. The results are shown in Table 4. UPS outperforms the non-metadata baseline, demonstrating the effect of incorporating metadata as a textual form of domain knowledge, which LLMs are able to understand. The results also suggest that feature concatenation is better than cross-modal attention, possibly because the latter is harder to optimize. We leave studying the optimal combination of metadata and PDE data as a future direction.
@@
S4: Other LLMs/VLMs To study whether UPS applies to other pretrained models, we further investigate Flan-T5 [47] and the vision language model CLIP [23]. In particular, for CLIP, we use its text model to encode the metadata and its vision model to process the PDE data. The results are reported in Table 4. Since these models are trained using the same datasets and optimizer configuration as RoBERTa, the results are not fully optimized. Nonetheless, their performance is competitive with existing baselines, and CLIP further outperforms RoBERTa on 3 tasks. This shows the compatibility of UPS with diverse pretrained backbones. A future direction is to study whether optimizing the training hyperparameters for each pretrained model—especially VLMs like CLIP that are trained for an additional vision modality—could improve downstream performance.
@@
S5: FNO Embedder & Target Sequence Length As discussed in Section 3, the channel $l$ of the FNO layers in the embedding network determines the sequence length of the PDE features that will be fed into the transformer layers. To study how this hyperparameter affects learning outcomes, we vary $l\in\{8,20,32\}$ and report the results in Table 4. In general, increasing $l$ improves the size and capacity of the embedding network, as well as the expressivity of the PDE features. This leads to lower prediction error. However, using too many parameters for the embedding network may result in a trade-off between effectiveness and efficiency. For instance, we also experimented with $l=64$ (Appendix D.3) and find that the longer sequence length leads to slight performance improvements but with much higher computational costs. Thus, we opt for $l=32$ in our main experiments.

### v4

@@ 3.2 Unified Architecture
Transformer models have demonstrated success in various domains like natural language (e.g., Touvron et al., 2023), vision (e.g., Dosovitskiy et al., 2021), and audio processing (e.g., Lu et al., 2023). In this work, we explore the potential of transformers for PDE solving. We break down the UPS architecture into 3 parts: an embedding network that transforms the unified representation into sequence features; the model body, consisting of the pretrained LLM layers; and a predictor that generates the prediction (Figure 1).
@@
We use FNO due to its strong empirical performance (Li et al., 2020a; Takamoto et al., 2022) and its ability to extract resolution-invariant features. As we consider maximum two-dimensional PDEs in this work, we use a series of 2D FNO layers with $l$ channels to obtain PDE features in $\mathbb{R}^{l\times n^{d}}$. Then, to map the FNO output to the LLM’s embedding dimension, we apply a pointwise convolution with input channel $n^{d}$, output channel $e$, kernel size $1$, stride $1$. This yields the desired sequential features $h_{\mathrm{PDE}}\in\mathbb{R}^{l\times e}$.
Since UPS is intended to handle diverse data from various generating sources, we leverage the PDE’s metadata in addition to the input dynamics. The motivation is that LLMs can use the textual information to better understand the context and characteristics of different PDEs. To implement this, we specify the metadata in the form “[PDE family][coefficients]” which is embedded into sequential features $h_{\mathrm{meta}}$ using the LLM’s tokenizer and text embedding layer. We then concatenate the meta features and the PDE features to get $h_{\mathrm{mix}}:=[h_{\mathrm{meta}},h_{\mathrm{PDE}}]$. Finally, we apply positional encoding and layer norm to $h_{\mathrm{mix}}$. This will be the input to the subsequent transformer layers.
@@
Linear Predictor Finally, we define a prediction head to transform the hidden state of the LLM body $\hat{h}$ to the predicted next step of the input $\hat{\bm{u}}_{t+1}^{s}(\bm{x})\in\mathbb{R}^{N\times n^{d}}$ (we predict all the physical quantities in the set $V$). This is achieved by averaging over the sequence dimension of $\hat{h}$ to get shape $\mathbb{R}^{e}$, applying a linear layer to map it to $\mathbb{R}^{Nn^{d}}$, and reshaping the results to obtain the final prediction $\hat{\bm{u}}_{t+1}^{s}(\bm{x})$. The linear predictor is shared for all PDEs.
@@ 4 Full Workflow and Training
Stage 1: Embedding Pretraining Intuitively, there is a modality gap between text data used to train general-purpose LLMs and PDEs. Previous work has also shown that directly fine-tuning pretrained LLMs on non-text inputs can result in suboptimal performance (Lu et al., 2022). To address this, Shen et al. (2023) introduced ORCA, which performs distribution matching before fine-tuning to enable cross-modal adaptation. That is, given a randomly initialized embedding network, we first pretrain it to minimize the distribution distance between the embedding network’s output—in our case $h_{\mathrm{mix}}$—and the text embeddings of a external reference NLP dataset, which we denote as $h_{\mathrm{LM}}$. This process makes the cross-modal distribution resemble the text distribution that the LLM is pretrained on. Following the ORCA work, we use the CoNLL-2003 dataset (Sang & Meulder, 2003) as the reference dataset for alignment.
@@
$\begin{split}{\mathcal{L}}_{\mathrm{align}}=\|\mu_{\mathcal{D}_{\mathrm{h_{mix}}}}-\mu_{\mathcal{D}_{\mathrm{h_{LM}}}}\|_{L_{2}}=\mathbb{E}_{\mathcal{D}_{\mathrm{h_{mix}}}}[k(a,a^{\prime})]-2\mathbb{E}_{\mathcal{D}_{\mathrm{h_{mix}}},\mathcal{D}_{\mathrm{h_{LM}}}}[k(a,b)]-\mathbb{E}_{\mathcal{D}_{\mathrm{h_{LM}}}}[k(b,b^{\prime})]\end{split}$ (2)
@@ 5 Experiments
Data We train and evaluate our method using PDEBench (Takamoto et al., 2022). For training, we combine 7 datasets from different PDE families: Burgers Equation (1D), Advection (1D), Diffusion-Sportion (1D), Shallow-Water (2D), compressible Navier-Stokes (1D and 2D), and incompressible Navier-Stokes (2D). We explicitly hold out two families, 1D and 2D Diffusion-Reaction, to evaluate the generalization ability of UPS. The dataset details can be found in Appendix A.1. We autoregressively generate the predictions and use the scale-independent normalized root mean squared error (nRMSE) as the evaluation metric, defined as follows:
@@
We preprocess all the PDEs by normalizing each dataset along the channel dimension to ensure the scale of $\bm{u}^{s}_{t}(\bm{x})$ across datasets is similar [^3] We standardize the data by subtracting the mean and dividing by the standard deviation to ensure training stability when using pretrained model weights. For loss computation, we apply an inverse transformation to the outputs to revert them to the original scale. Although data normalization may affect non-linear equations, it is essential to prevent loss explosion during fine-tuning. We leave exploring alternative methods to minimize distortion in non-linear dynamics as future work..
Baselines We compare against two sets of baselines: (i) single-family models trained on individual PDE datasets, including the widely used U-Net (Ronneberger et al., 2015), FNO (Li et al., 2020b), the improved version FFNO (Tran et al., 2023), the transformer-based GNOT (Hao et al., 2023b) and OFormer (Li et al., 2023), as well as the cross-modal ORCA (Shen et al., 2023); (ii) unified models trained on multiple datasets, including MPP (McCabe et al., 2023), DPOT (Hao et al., 2024), and a unified FNO trained using data transformed by our unified representation scheme. We note that MPP and DPOT focus on 2D PDEs and are pretrained on 2D Navier-Stokes, Shallow-Water, and Diffusion-Reaction from PDEBench. Subramanian et al. (2023) is not included as a baseline because its models are pretrained on different PDE families (e.g., Poisson’s and Helmholtz equations) not present in PDEBench.
Table 1: nRMSEs (lower is better) for in-distribution PDEBench families, with baseline results taken from Takamoto et al. (2022); Shen et al. (2023); McCabe et al. (2023); Hao et al. (2024). ‘-’ means that the result is not available. On all datasets, UPS with RoBERTa-Base (UPS-B) achieves the lowest nRMSEs among all smaller models and UPS with RoBERTa-Large (UPS-L) achieves the lowest nRMSEs among all large models. Numbers are bolded for each model size group.
@@
Implementation Details As noted in Section 3, UPS is compatible with any pretrained LLM. We present our main results using RoBERTa (Liu et al., 2019) and study other backbones in ablation studies (Table 4). We set the embedding FNO channel $l$ to 32. Since the resolution of the 2D datasets in PDEBench is 128, we set the model resolution $n$ to 128 and downsample datasets with higher resolutions. All of our experiments can be run on a single NVIDIA A6000 GPU. See Appendix A.2 for training details. Due to computational constraints, results are based on a single run per network configuration.
Since MPP and DPOT focus on 2D problems and use a different set of pretraining datasets from ours, we train a 2D-only UPS on all 2D datasets in PDEBench to provide a more direct comparison. The results are shown in Appendix Table 8. Notably, while 2D UPS is still trained with less data (since the other methods use additional datasets like PDEArena (Gupta & Brandstetter, 2022)), our method ranks first on 4 of 8 datasets, outperforming DPOT on 5 of 8 datasets and outperforming MPP on 3 of 4 datasets.
@@ 5.1 Achieving State-of-the-Art Results on PDEBench with Compute Efficiency
In addition to prediction errors in Table 1, we visualize some of the UPS outputs in Appendix A.4 and show that it is indeed able to capture the key features and dynamics of different PDE families. For efficiency metrics, we report the training compute requirement, FLOPs, and inference time for UPS in Appendix A.2.2. Compared to existing work, our method has lower FLOPs and shorter inference time. This shows that our method can be deployed in practical environments where both computational efficiency and speed are critical.
@@ 5.2 Generalizing to Unseen PDEs with Data Efficiency
Table 2: Zero- and few-shot transfer performance of UPS on unseen PDE families and coefficients. Our few-shot results are competitive with baselines trained with more data. UPS-B refers to UPS with RoBERTa-Base.
Unseen PDE Families Unseen Coefficients
# Samples Model 1D Diff-React 2D Diff-React 1D Burgers 2D Navier-Stokes
$\nu=1.0$ $M=1$, $\eta=\zeta=0.1$
UPS-B 0.0557 1.0593 0.0566 0.103
0 FNO 0.1839 1.2 1.0342 1.4302
ORCA 0.1818 1.0812 1.6316 1.6399
UPS-B 0.0107 0.3327 0.0134 0.0809
10 FNO 0.1698 0.8193 0.67 0.567
ORCA 0.1004 0.5376 0.4829 0.1623
UPS-B 0.0034 0.2508 0.0022 0.0543
100 FNO 0.0037 0.1869 0.0123 0.3962
ORCA 0.0051 0.1362 0.027 0.0898
UPS-B 0.0003 0.041 0.0008 0.0191
9K (Full) FNO 0.0014 0.12 0.0031 0.098
ORCA 0.0034 0.082 0.012 0.0287
@@
Unseen PDE Families As mentioned earlier, we hold out the Diffusion-Reaction equations from developing UPS. We first directly evaluate UPS on these two tasks and report the zero-shot transfer performance. Then, we study few-shot transfer by randomly sampling $k\in\{10,100\}$ trajectories from the training sets of the held-out tasks and use them to fine-tune UPS. Lastly, we report the fine-tuning results with the full training dataset. The results are shown in Table 2. As the number of adaptation samples increases, the prediction error decreases. Notably, the 100-shot result of UPS on 1D datasets is better than the baselines trained on 9,000 data, i.e., we use 90x less data to match the performance of single-family operators. This makes UPS useful for low-resource PDE problems where data collection is costly and training models from scratch is challenging. On 2D Diffusion-Reaction, we are slightly worse than pretrained MPP (0.0292) and DPOT (0.0106) since this dataset is considered as in-distrubution for MPP and DPOT.
Unseen Coefficients UPS also generalizes to PDEs in the same families as the training data but with different coefficients. We verify this by adapting UPS to Burgers Equation with $\nu=1.0$ (the model is trained on $\nu=0.001$) and compressible Navier-Stokes with $M=1$, $\eta=\zeta=0.1$ (the model is trained on $M=\eta=\zeta=0.1$). The last two columns in Table 2 shows that while our zero-shot performance is already competitive, the performance after further adaptation outperforms most considered baselines.
Table 3: UPS with resolution 128 has an nRMSE of 0.0033 for Advection and 0.0931 for incompressible Navier-Stokes. We directly test UPS on higher resolutions.
@@
Advection (nRMSE) 0.0057 0.0064 0.0068
Incomp Navier-Stokes (nRMSE) 0.119 0.126 -
@@
We report the resolution generalization performance for 1D Advection Equation and 2D incompressible Navier-Stokes in Table 3. Although the nRMSEs for both PDEs slightly increase compared to the nRMSE for the training resolution, they outperform all baselines in Table 1. Since the numbers are similar across columns, UPS generalizes to higher resolutions in a zero-shot manner.
@@ 5.3 Ablation Studies
Table 4: Results for the ablation studies. For each set of experiments, only the specified settings are different; all the other hyperparameters and training configurations are the same. Overall, the full UPS-Base workflow (first row for every study) most effectively leverages the pretrained knowledge of LLMs and obtains the best results.
@@
S3 Concatenate Pretrained Text 0.0027 0.0399 0.0009 0.0056 0.0016 0.0153 0.0931
Cross-Attention Pretrained Text 0.003 0.0420 0.0009 0.0065 0.0023 0.0189 0.1082
Concatenate One-Hot 0.0029 0.0447 0.0011 0.006 0.0018 0.0198 0.095
Concatenate Learned Embeddings 0.0041 0.0474 0.0014 0.0119 0.0036 0.0295 0.1103
No Meta Information 0.0122 0.0453 0.001 0.0091 0.0026 0.0238 0.1171
@@
S2: Cross-Modal Alignment We also test the importance of the two objectives used in stage 1, i.e., alignment loss with MMD, and task loss with nRMSE. We study three settings: (i) using only ${\mathcal{L}}_{\mathrm{align}}$ for stage 1 as in Shen et al. (2023); (ii) using only ${\mathcal{L}}_{\mathrm{task}}$ for stage 1; and (iii) removing stage 1 from our workflow entirely. As shown in Table 4, while removing any objective reduces the performance across all datasets, removing the task loss has a more significant negative effect. Meanwhile, removing the entire stage of embedding pretraining hurts prediction accuracy. This shows that simply fine-tuning the LLM without considering the modality gap or learning to extract PDE features is ineffective.
S3: Incorporating Text-Form Metadata UPS leverages the PDE’s metadata by combining its text embeddings with the learned PDE embeddings. To study whether incorporating such metadata is helpful and identify an optimal approach, we compare our workflow with two alternatives: (i) we do not use metadata, so $h_{\mathrm{mix}}:=h_{\mathrm{PDE}}$; (ii) we use metadata, but instead of concatenating features from two modalities, we apply a cross-attention mechanism: $h_{\mathrm{mix}}:=\mathrm{softmax}(\frac{QK^{T}}{\sqrt{e}})V$, where $Q=W_{Q}h_{\mathrm{PDE}}$, $K=W_{K}h_{\mathrm{meta}}$, and $V=W_{V}h_{\mathrm{meta}}$. To further investigate whether the pretrained text embeddings contribute beyond merely labeling the PDE type, we also study alternative embedding strategies that do not leverage language pretraining: (iii) we replace the pretrained text embeddings of the PDE meta information with one-hot encoded vectors representing each PDE type. This setting serve as a baseline to assess the impact of merely labeling the PDE types without any semantic understanding; (iv) we also test a setting where PDE types were embedded using a randomly initialized embedding layer that was trained from scratch along with the rest of the network, i.e., each new token represents a PDE family.
@@
The results are shown in Table 4. UPS outperforms the non-metadata baseline, demonstrating the effect of incorporating metadata as a textual form of domain knowledge, which LLMs are able to understand. The results also suggest that feature concatenation is better than cross-modal attention, possibly because the latter is harder to optimize. Lastly, the setting utilizing pretrained text embeddings consistently outperforms the one-hot and embedding-from-scratch settings. In terms of learning dynamics, we also observe that using pretrained embeddings demonstrated faster convergence compared to the alternative strategies. This suggests that the pretrained semantic knowledge in the LLM indeed contributes to processing the PDE data, not just in labeling PDE types but might also in understanding the underlying physical phenomena. However, we leave studying the the exact mechanism of cross-modal transfer and the optimal combination of metadata and PDE data as a future direction.
@@
S4: Other LLMs/VLMs To study whether UPS applies to other pretrained models, we further investigate Flan-T5 (Chung et al., 2022) and the vision language model CLIP (Radford et al., 2021). In particular, for CLIP, we use its text model to encode the metadata and its vision model to process the PDE data. The results are reported in Table 4. Since these models are trained using the same datasets and optimizer configuration as RoBERTa, the results are not fully optimized. Nonetheless, their performance is competitive with existing baselines, and CLIP further outperforms RoBERTa on 3 tasks. This shows the compatibility of UPS with diverse pretrained backbones. A future direction is to study whether optimizing the training hyperparameters for each pretrained model—especially VLMs like CLIP that are trained for an additional vision modality—could improve downstream performance.
@@
S5: FNO Embedder & Target Sequence Length As discussed in Section 3, the channel $l$ of the FNO layers in the embedding network determines the sequence length of the PDE features that will be fed into the transformer layers. To study how this hyperparameter affects learning outcomes, we vary $l\in\{8,20,32\}$ and report the results in Table 4. In general, increasing $l$ improves the size and capacity of the embedding network, as well as the expressivity of the PDE features. This leads to lower prediction error. However, using too many parameters for the embedding network may result in a trade-off between effectiveness and efficiency. For instance, we also experimented with $l=64$ (Appendix A.3.3) and find that the longer sequence length leads to slight performance improvements but with much higher computational costs. Thus, we opt for $l=32$ in our main experiments.